OpenStudy (anonymous):

find Derivative of 2^(-x) ?

OpenStudy (anonymous):

do you know what the derivative of \(b^x\) for arbitrary base \(b\) ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\frac{d}{dx}[b^x]=b^x\ln(b)\]

OpenStudy (anonymous):

so \[\frac{d}{dx}2^{-x}=-2^{-x}\ln(2)\] by the chain rule

OpenStudy (anonymous):

there should be e right ?